# What Matters in Transformers? Not All Attention is Needed

### v2

@@ Abstract
Scaling Transformer-based large language models (LLMs) has demonstrated promising performance across various tasks. However, this scaling also introduces redundant structures, posing challenges for real-world deployment. Despite some recognition of redundancy in LLMs, the variability of redundancy across different structures, such as MLP and Attention layers, is under-explored. In this work, we investigate the varying redundancy across different modules within Transformers, including Blocks, MLP, and Attention layers, using a similarity-based metric. This metric operates on the premise that redundant structures produce outputs highly similar to their inputs. Surprisingly, while attention layers are essential for transformers and distinguish them from other mainstream architectures, we found that a large proportion of attention layers exhibit excessively high similarity and can be safely pruned without degrading performance, leading to reduced memory and computation costs. Additionally, we further propose a method that jointly drops Attention and MLP layers, achieving improved performance and dropping ratios. Extensive experiments demonstrate the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. Our findings provide valuable insights for future network architecture design. The code will be released at: https://github.com/Shwai-He/LLM-Drop.
@@ 1 Introduction
Given that a Transformer block comprises an MLP layer and an Attention layer, we examine the effects of dropping these layers separately. Our findings reveal that while dropping MLP layers negatively impacts performance, dropping Attention layers, i.e., the core of Transformer architectures which distinguish it from other mainstream architectures (e.g., RWKV Peng et al. (2023) and Mamba Gu and Dao (2024)), can be done extensively without degrading model performance. For example, dropping 50% of the Attention layers in Llama-2-70B Touvron et al. (2023) results in comparable performance to the original model, indicating a high level of redundancy in these layers.
@@ 2 Related Works
Model Compression. LLMs can be compressed to promote their efficiency in memory and computation. Quantization Frantar et al. (2023); Lin et al. (2024) and Pruning Sun et al. (2023); Frantar and Alistarh (2023) are the most widely used techniques to compress LLMs. Specifically, quantization transforms the data type into low-bit but remains potentially redundant architecture and parameters. Pruning can be categorized into unstructured pruning and structured pruning, where the former does not apply to hardware while the latter degrades the performance significantly. Compared to pruning, our methods, i.e., Block Drop and Layer Drop, not only make the model architectures more compact but also maintain the performance. On the other hand, Block Drop and Layer Drop are orthogonal to quantization, and their integration with quantization significantly enhances the efficiency.
@@ 3.1 Similarity-based Drop
$\boldsymbol{S}=1-Cosine(\boldsymbol{X},\boldsymbol{Y}),$ (1)
@@ 3.2 Block Drop
$\boldsymbol{S}^{l}_{{B}}=1-Cosine(\boldsymbol{X}^{l}_{B},\boldsymbol{Y}^{l}_{B}),$ (2)
where $\boldsymbol{X}^{l}_{B}$ and $\boldsymbol{Y}^{l}_{B}$ denote the input and output of the $l$-th block, respectively. Since the similarity scores are computed locally, we can offload irrelevant modules to save memory. Specifically, we iteratively compute the importance score of each block from shallow to deep, and then drop the blocks with the lowest scores.
@@
where $\boldsymbol{Y}^{l}_{{M}}$ involves two parts, and the output of the MLP constitutes only a portion of it. Therefore, directly utilizing the similarity metric on the inputs and outputs of the MLP leads to inaccurate measurements. To capture the overall inputs and outputs accurately, we consider the MLP layers together with the subsequent LayerNorm as a single unit for dropping. This approach ensures a more accurate measurement. The importance metrics are then as follows:
$\boldsymbol{S}^{l}_{{M}}=1-Cosine(\boldsymbol{X}^{l}_{M},\boldsymbol{Y}^{l}_{M}).$ (4)
Note that we consider MLP layers together with their corresponding LayerNorm layers. Therefore, MLP Drop removes both the insignificant MLP layers and the associated LayerNorm layers.
Attention Drop. Similar to MLP Drop, Attention Drop also considers the residual connection, as the outputs of attention layers contribute only partially to the overall outputs:
@@
where $\boldsymbol{X}^{l}_{{A}}$ is the inputs of the corresponding LayerNorm layers and $\boldsymbol{Y}^{l}_{{A}}$ overall outputs that involves residual connections. Like MLP Drop that takes attention layers and associated LayerNorm layers together, the importance score for Attention Drop is:
$\boldsymbol{S}^{l}_{{A}}=1-Cosine(\boldsymbol{X}^{l}_{A},\boldsymbol{Y}^{l}_{A}).$ (6)
Similar Block Drop, the computation of importance scores of Layer Drop is conducted layer-wisely from shallow layers to deep layers. All Drop methods are conducted with a calibration dataset $\mathcal{D}$ (e.g., C4 Raffel et al. (2020)) in a one-shot manner that computes the importance scores once and then drops redundant layers, instead of using resource-consuming multi-turn iterations. The effectiveness of One-Shot Drop has been verified in Appendix B. On the other hand, other than solely dropping Attention or MLP layers individually, we propose Joint Layer Drop that considers the combined importance scores across these layers to optimize the dropping process. The algorithm for Joint Layer Drop is presented in Algorithm 1.
@@
Figure 4: Visualization of Importance Scores in Checkpoints during the Pre-training Process of MAP-Neo-7B, where lighter areas represent low importance scores (i.e., high similarity scores). We present both (1) Early Stage (the first 12 released checkpoints) and (2) Entire Process (checkpoints for every 500B trained tokens). We independently visualize the importance scores at the module index 0, since they are significantly higher.
@@ 4.1 Main Results
We next delve into removing fine-grained modules in Blocks, i.e., MLP and Attention Layers. Similar to Block Drop, dropping MLP layers also leads to a substantial performance decline, e.g., 6.3% on Llama-2-13B and 6.9% on Mistral-7B when dropping 8 layers, which is unacceptable for practical usage.
Surprisingly, dropping attention layers demonstrates high effectiveness, as models with 8 dropped layers still maintain most of their performance, e.g., both Llama-3-13B and Mistral-7B maintain more than 99% of the original performance after dropping 8 attention layers. Although attention layers are the hallmarks of transformers, dropping such layers does not degrade the performance.
@@ 4.2 Detailed Analysis
Figure 6: Accuracy Curves of Dropping Different Target Modules, where we consider dropping single types of modules and Joint dropping (Attn + MLP). In the Joint’s line, $\bigstar$ represents the step where the MLP is dropped, while the $\blacksquare$ represents the step where the Attention is dropped.
Block Drop and Layer Drop are Data Robust. In Figure 7 and Figure 11, we note that feature similarity remains relatively stable across different Attention layers as the sample size increases (i.e., from 4 to 1024), indicating that Block Drop and Layer Drop maintain consistency regardless of sample quantity. This confirms that using $256$ samples suffice for computing similarity, which is adopted for all our experiments. Similarly, Figure 8 and Figure 12 show that varying the calibration datasets, from the pretraining dataset (e.g., C4 Raffel et al. (2020)) to instruction tuning datasets (e.g., CodeAlpaca-20k [^1] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k, MathInstruct Yue et al. (2024) and LIMA Zhou et al. (2024)), does not significantly alter feature similarity. Despite some differences, the modules dropped remain the same. This demonstrates the resilience of Block Drop and Layer Drop to variations in calibration datasets.
@@
Similar to the findings in smaller models, Llama-2-70B also showcases sensitivity to Block Drop and Layer Drop, where dropping only 20% of blocks or MLP layers leads to a significant performance drop. However, it is noted that larger models are more robust than smaller models: Llama-2-70B dropped with 8 MLP layers only encounters 0.6% performance drop on average, compared to 3.9% of Llama-2-13B when dropping 4 MLP layers. Table 3 also reveals a similar trend on Llama-3.
On the other hand, Attention Drop performs well on Llama-2-70B. Specifically, when dropping 32 out of 80 attention layers, the performance of Llama-2-70B decreases by only 0.1% on average. Conversely, on Llama-2-13B, dropping the same proportion of attention layers leads to a noticeable performance drop.
Joint Layer Drop Enhances the Performance. In this part, we assemble two strategies of Layer Drop, i.e., Attention Drop and MLP Drop. Specifically, we first obtain the importance scores ($\boldsymbol{S}_{{A}}$ and $\boldsymbol{S}_{{M}}$) for Attention layers and MLP Layers, respectively. Then, we concatenate the importance scores ($\boldsymbol{S}=[\boldsymbol{S}_{{A}},\boldsymbol{S}_{{M}}]$), and drop the layers with the lowest scores.
@@ Appendix B Additional Experimental Results
To evaluate Attention Drop on more complex technical tasks, we evaluated Llama-2-7B and and Mistral-7B, and two corresponding instruction fine-tuned models, meta-math/MetaMath-7B-V1.0 and meta-math/MetaMath-Mistral-7B Yu et al. (2023). The results in Figure 9 indicate that, except for Llama-2-7B-Math, which is meta-math/MetaMath-7B-V1.0, all the models do not experience significant performance degradation when dropping fewer than 8 Attention layers. We speculate that this is because Llama-2-7B-Math is initialized with Llama-2-7B and undergoes instruction fine-tuning to improve its mathematical ability. Llama-2-7B-Base exhibits poor performance in mathematics, and the ability obtained solely through fine-tuning appears to be superficial. Therefore, when dropping Attention layers, Llama-2-7B-Math’s ability rapidly deteriorates.
@@
Iterative and One-Shot are the most common settings for model compression, where the former prunes step by step while one-shot prune the model once. Iterative pruning may causes more accurate measurements but tend to be more time-consuming. In Figure 10, we empirically compare Iterative Dropping and One-Shot Dropping. For Iterative Dropping, we consider dropping one by one iteratively.

### v3

@@ Abstract
Scaling Transformer-based large language models (LLMs) has demonstrated promising performance across various tasks. However, it also introduces redundant structures, posing challenges for real-world deployment. Despite some recognition of redundancy in LLMs, the variability of redundancy across different modules, such as MLP and Attention layers, is under-explored. In this work, we investigate the varying redundancy across different modules within Transformers, including Blocks, MLP, and Attention layers, using a similarity-based metric. This metric operates on the premise that redundant structures produce outputs highly similar to their inputs. Surprisingly, while attention layers are essential for transformers and distinguish them from other mainstream architectures, we found that a large proportion of attention layers exhibit excessively high similarity and can be safely pruned without degrading performance, leading to reduced memory and computation costs. Additionally, we further propose a method that jointly drops Attention and MLP layers, achieving improved performance and dropping ratios. Extensive experiments demonstrate the effectiveness of our methods, e.g., Llama-3-70B maintains comparable performance even after pruning half of the attention layers. Our findings provide valuable insights for future network architecture design. The code is released at: https://github.com/Shwai-He/LLM-Drop.
@@ 1 Introduction
Given that a Transformer block comprises an MLP layer and an Attention layer Vaswani et al. (2023), we examine the effects of dropping these layers separately. Our findings reveal that while dropping MLP layers negatively impacts performance, dropping Attention layers, i.e., the core of Transformer architectures which distinguish it from other mainstream architectures (e.g., RWKV Peng et al. (2023) and Mamba Gu and Dao (2024)), can be done extensively without degrading model performance. For example, dropping 50% of the Attention layers in Llama-2-70B Touvron et al. (2023) results in comparable performance to the original model, indicating a high level of redundancy in these layers.
@@ 2 Related Works
Model Compression. LLMs can be compressed to promote their efficiency in memory and computation. Quantization Frantar et al. (2023); Lin et al. (2024) and Pruning Sun et al. (2023); Frantar and Alistarh (2023) are the most widely used techniques to compress LLMs. Specifically, quantization transforms the data type into low-bit but remains potentially redundant architecture and parameters. Pruning can be categorized into unstructured pruning Kusupati et al. (2020); Sanh et al. (2020) and structured pruning Zhuang et al. (2020); Kwon et al. (2020), where the former does not apply to hardware while the latter degrades the performance significantly. Compared to pruning, our methods, i.e., Block Drop and Layer Drop, not only make the model architectures more compact but also maintain the performance. On the other hand, Block Drop and Layer Drop are orthogonal to quantization, and their integration with quantization significantly enhances the efficiency.
@@ 3.1 Similarity-based Drop
$\boldsymbol{S}=1-\text{Cosine}(\boldsymbol{X},\boldsymbol{Y}),$ (1)
@@ 3.2 Block Drop
$\boldsymbol{S}^{l}_{{B}}=1-\text{Cosine}(\boldsymbol{X}^{l}_{B},\boldsymbol{Y}^{l}_{B}),$ (2)
where $\boldsymbol{X}^{l}_{B}$ and $\boldsymbol{Y}^{l}_{B}$ denote the input and output of the $l$-th block, respectively. Since the similarity scores are computed locally, we can offload irrelevant modules to save memory. Specifically, we sequentially compute the importance score of each block from shallow to deep, and then drop the blocks with the lowest scores.
@@
where $\boldsymbol{Y}^{l}_{{M}}$ involves two parts, and the output of the MLP layer constitutes only a portion of it. Therefore, directly utilizing the similarity metric on the inputs and outputs of an MLP layer leads to inaccurate measurements. To capture the overall inputs and outputs accurately, we consider an MLP layer together with its subsequent LayerNorm as a single unit for dropping. This approach ensures a more accurate measurement. The importance metrics are then as follows:
$\boldsymbol{S}^{l}_{{M}}=1-\text{Cosine}(\boldsymbol{X}^{l}_{M},\boldsymbol{Y}^{l}_{M}).$ (4)
Note that we consider MLP layers together with their corresponding LayerNorm layers. Therefore, MLP Drop removes both the insignificant MLP layers and their associated LayerNorm layers.
Attention Drop. Similarly, Attention Drop also considers the residual connection, as the outputs of attention layers contribute only partially to the overall outputs:
@@
where $\boldsymbol{X}^{l}_{{A}}$ is the inputs of the corresponding LayerNorm layers and $\boldsymbol{Y}^{l}_{{A}}$ overall outputs that involves residual connections. Like MLP Drop associates LayerNorm layers together, the importance score for Attention Drop is:
$\boldsymbol{S}^{l}_{{A}}=1-\text{Cosine}(\boldsymbol{X}^{l}_{A},\boldsymbol{Y}^{l}_{A}).$ (6)
Similar Block Drop, the computation of importance scores of Layer Drop is conducted layer-wisely from shallow layers to deep layers. All drop methods are conducted with a calibration dataset $\mathcal{D}$ (e.g., C4 Raffel et al. (2020)) in a one-shot manner that computes the importance scores once and then drops all redundant layers, instead of using resource-consuming multi-turn iterations. The effectiveness of One-Shot Drop has been verified in Appendix B. On the other hand, other than solely dropping Attention or MLP layers individually, we propose Joint Layer Drop that considers the combined importance scores across Attention and MLP to optimize the dropping process. The algorithm for Joint Layer Drop is presented in Algorithm 1.
@@
Figure 4: Visualization of Importance Scores in Checkpoints during the Pre-training Process of MAP-Neo-7B, where light areas represent low importance scores (i.e., high similarity scores). We present both (1) Early Stage (the first 12 released checkpoints) and (2) Entire Process (checkpoints for every 500B trained tokens). We independently visualize the importance scores at the module index 0, since they are significantly higher.
@@ 4.1 Main Results
We next delve into the removal of more fine-grained modules in Blocks, i.e., MLP and Attention Layers. Similar to Block Drop, MLP Drop also leads to a substantial performance decline, e.g., 6.3% on Llama-2-13B and 6.9% on Mistral-7B when dropping 8 layers, which is unacceptable for practical usage.
Conversely, Attention Drop demonstrates surprising effectiveness, as models with 8 dropped layers still maintain most of their performance, e.g., both Llama-3-13B and Mistral-7B maintain more than 99% of the original performance. Although attention layers are the hallmarks of transformers, dropping such layers does not degrade the performance significantly.
@@ 4.2 Detailed Analysis
Figure 6: Accuracy Curves of Dropping Different Target Modules, where we consider dropping single types of modules and Joint Drop (Attn + MLP). In the line of Joint Drop, $\bigstar$ represents the step where the MLP is dropped, while the $\blacksquare$ represents the step where the Attention is dropped.
Block Drop and Layer Drop are Data Robust. In Figure 7 and Figure 11, we note that feature similarity remains relatively stable across different modules of Mistral-7B as the sample size increases (i.e., from 4 to 1024), indicating that Block Drop and Layer Drop maintain consistency regardless of sample quantity. This confirms that using $256$ samples suffice for computing similarity, which is adopted for all our experiments. Similarly, Figure 8 and Figure 12 show that varying the calibration datasets, from the pretraining dataset (e.g., C4 Raffel et al. (2020)) to instruction tuning datasets (e.g., CodeAlpaca-20k [^1] https://huggingface.co/datasets/sahil2801/CodeAlpaca-20k, MathInstruct Yue et al. (2024) and LIMA Zhou et al. (2024)), does not significantly alter feature similarity. Despite some differences, the modules dropped remain the same. This demonstrates the resilience of Block Drop and Layer Drop to variations in calibration datasets.
@@
Similar to the findings in smaller models, Llama-2-70B also showcases sensitivity to Block Drop and Layer Drop, where dropping only 20% of blocks or MLP layers leads to a significant performance drop. Differently, it is noted that larger models are more robust than smaller models: Llama-2-70B dropped with 8 MLP layers only encounters 0.6% performance drop on average, compared to 3.9% of Llama-2-13B when dropping 4 MLP layers. Table 3 also reveals a similar trend on Llama-3.
On the other hand, Attention Drop performs well on Llama-2-70B. Specifically, when dropping 32 out of 80 attention layers, the performance of Llama-2-70B decreases by only 0.1% on average. Conversely, on Llama-2-13B, dropping the same proportion of attention layers leads to a more noticeable performance drop.
Joint Layer Drop Enhances the Performance. In this part, we assemble two strategies of Layer Drop, i.e., Attention Drop and MLP Drop. Specifically, we first obtain the importance scores ($\boldsymbol{S}_{{A}}$ and $\boldsymbol{S}_{{M}}$) for Attention layers and MLP Layers, respectively. Then, we concatenate the importance scores ($\boldsymbol{S}=\{\boldsymbol{S}_{{A}},\boldsymbol{S}_{{M}}\}$), and drop the layers with the lowest scores.
@@ Appendix B Additional Experimental Results
To evaluate Attention Drop on more complex technical tasks, we evaluated Llama-2-7B and and Mistral-7B, and two corresponding instruction fine-tuned models, meta-math/MetaMath-7B-V1.0 and meta-math/MetaMath-Mistral-7B Yu et al. (2023). The results in Figure 9 indicate that, except for Llama-2-7B-Math, which is meta-math/MetaMath-7B-V1.0, all the models do not experience significant performance degradation when dropping fewer than 8 Attention layers. We speculate that this is because Llama-2-7B-Math is initialized with Llama-2-7B and undergoes instruction fine-tuning to improve its mathematical ability. Instead, Llama-2-7B-Base exhibits poor performance in mathematics, and the ability obtained solely through fine-tuning appears to be superficial. Therefore, when dropping Attention layers, Llama-2-7B-Math’s ability rapidly deteriorates.
@@
Iterative and One-Shot are the most common settings for model compression, where the former prunes step by step and the latter prunes the model once. Iterative pruning may causes more accurate measurements but tend to be more time-consuming. In Figure 10, we empirically compare Iterative Dropping and One-Shot Dropping. For Iterative Dropping, we consider dropping one by one iteratively.